# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v3

@@ Abstract
The aim of this paper is to bridge noncommutative geometry with classical harmonic analysis on Banach spaces, focusing primarily on both classical and noncommutative $\mathrm{L}^{p}$ spaces. Introducing a notion of Banach Fredholm module, we define new abelian groups, $\mathrm{K}^{0}(\mathcal{A},\mathscr{B})$ and $\mathrm{K}^{1}(\mathcal{A},\mathscr{B})$, of $\mathrm{K}$-homology associated with an algebra $\mathcal{A}$ and a suitable class $\mathscr{B}$ of Banach spaces, such as the class of $\mathrm{L}^{p}$-spaces. We establish index pairings of these groups with the $\mathrm{K}$-theory groups of the algebra $\mathcal{A}$. Subsequently, by considering (noncommutative) Hardy spaces, we uncover the natural emergence of Hilbert transforms, leading to Banach Fredholm modules and culminating in index theorems. Moreover, by associating each reasonable sub-Markovian semigroup of operators with a ¡¡Banach noncommutative manifold¿¿, we explain how this leads to (possibly kernel-degenerate) Banach Fredholm modules, thereby revealing the role of vectorial Riesz transforms in this context. Overall, our approach significantly integrates the analysis of operators on $\mathrm{L}^{p}$-spaces into the expansive framework of noncommutative geometry, offering new perspectives.
@@ 1 Introduction
A fundamental concept in this framework is that of a Fredholm module. If $\mathcal{A}$ is unital, a Fredholm module over $\mathcal{A}$ is a bounded operator $F\colon H\to H$ satisfying the three relations $^{2}=\mathrm{Id}_{H}, F=F$$F^{*}$ and $F,\pi(a)]=0$ for any $a\in\mathcal{A}$ modulo a compact operator. The last equality is a form of ¡¡pseudolocality¿¿. This notion of Fredholm module is a ¡¡normalized¿¿ variant of the notion of Atiyah cycle1110. An Atiyah cycle on $\mathcal{X}$ is a Fredholm operator $T\colon H_{1}\to H_{2}$ in Hilbert spaces $H_{1}$ and $H_{2}$ equipped with the structure of *-modules over the algebra $\mathrm{C}(\mathcal{X})$ such that the commutator $[T,f]$ is a compact operator for any function $f\in\mathrm{C}(\mathcal{X})$. (or generalized elliptic operator) introduced in the paper [Ati70] (see also [BDF77, Definition 7.6 p. 317], [Dou86, Definition 6 p. 165] and [NSSS06, Definition 10.9 p. 274]). If $D$ is the Dirac operator on a compact Riemannian spin manifold $M$, the Fredholm module $\operatorname{\mathrm{sgn}}D$ canonically associated to $M$ encodes the conformal structure of the manifold, as discussed in [Bar07, Theorem 3.1 p. 388]. These operators enable the definition of $\mathrm{K}$-homology groups $\mathrm{K}^{0}(\mathcal{A})$ and $\mathrm{K}^{1}(\mathcal{A})$, whose elements are equivalence classes of Fredholm modules, defined under appropriate notions of unitary equivalence and homotopy equivalence, and a natural operation of addition. The distinction between the two groups lies in the fact that, in the case of $\mathrm{K}^{0}(\mathcal{A})$, only even Fredholm modules are considered, provided an appropriate grading is introduced. These groups are linked to the $\mathrm{K}$-theory groups $\mathrm{K}_{0}(\mathcal{A})$ and $\mathrm{K}_{1}(\mathcal{A})$ of the algebra $\mathcal{A}$ through pairings
@@
of $a$. In the case $F=\mathrm{i}\mathcal{H}$, we have
(1.6) $\big([F,M_{a}]f\big)(x)=\operatorname{p.v.}\int_{-\infty}^{\infty}\frac{a(x)-a(y)}{x-y}f(y)\mathop{}\mathopen{}\mathrm{d}y.$
@@
by integrating over the boundary of a larger bisector $\Sigma^{\mathrm{bi}}_{\nu}$ using the resolvent operator $R(z,D)\overset{\mathrm{def}}{=}(z-D)^{-1}$, where $\omega<\nu<\theta$. Note that the boundedness of such a functional calculus is not free, contrary to the case of the functional calculus of a selfadjoint operator. We refer to our paper [Arh24c] for concrete examples of operator with such a bounded functional calculus, where we use a notion of curvature for obtaining it. Using the function $\operatorname{\mathrm{sgn}}$ defined by $\operatorname{\mathrm{sgn}}(z)\overset{\mathrm{def}}{=}1_{\Sigma_{\theta}}(z)-1_{-\Sigma_{\theta}}(z)$, we show as the hilbertian case that $\operatorname{\mathrm{sign}}D\colon X\to X$ is a Fredholm module. This notion of functional calculus was popularized in the paper [AKM06], which contains a (second) solution to famous Kato’s square root problem solved in [AHLMT02] and in [AKM06] (see also [HLM02] and [Tch01]).
@@ 2 Preliminaries
where $(\varepsilon_{k})_{k\geqslant 1}$ is a sequence of independent Rademacher variables on some probability space $\Omega$. As explained in [HvNVW18, Remark 8.1.2 p. 165], this property is independent of $p$. By [HvNVW18, Theorem 8.1.3 p. 166], an $R$-bounded subset of operators is bounded. Actually, the notion of $R$-boundedness can be viewed as a natural and powerful substitute for the usual notion of boundedness, particularly useful when extending results that hold on Hilbert spaces to the more general setting of Banach spaces.
@@
Let $\mathrm{H}^{\infty}_{0}(\Sigma_{\theta})$ be the subalgebra of bounded analytic functions $f\colon\Sigma_{\theta}\to\mathrm{C}$ for which there exist $s,c>0$ such that
@@
The following result is [HvNVW18, Proposition 10.6.2 p. 448].
@@
Let $\omega\in(0,\frac{\pi}{2})$. If $A$ is a bisectorial operator of type $\omega$ then $A^{2}$ is sectorial of type $2\omega$.
@@
We refer to the books [Bl98], [GVF01], [RLL00] and [WeO93] for more information. Two idempotents $e,f$ in a $\mathbb{C}$-algebra $\mathcal{A}$ are Murray von Neumann equivalent, written $e\sim f$, if there are $v,w\in S$ with $vw=e$ and $wv=f$. Let $\mathcal{A}$ be a unital $\mathbb{C}$-algebra, and let $\mathrm{V}(\mathcal{A})$ denote the collection of all Murray-von Neumann equivalence classes of idempotents in $\mathrm{M}_{\infty}(\mathcal{A})$. Equip $\mathrm{V}(\mathcal{A})$ with the addition operation defined by
$e]+[f]=\begin{bmatrix}e&0\\ 0&f\\ \end{bmatrix}$
Then $\mathrm{V}(\mathcal{A})$ becomes a commutative monoid. Let $\mathcal{A}$ be a unital complex algebra. Then idempotents $e,f\in\mathrm{M}_{\infty}(\mathcal{A})$ are Murray-von Neumann equivalent if and only if there is an invertible element $u$ in some $\mathrm{M}_{n}(\mathcal{A})$ such that $ueu^{-1}=f$.
@@
Let $A$ be a unital $\mathbb{C}$-algebra. The group $\mathrm{K}_{0}(\mathcal{A})$ is the Grothendieck group of $\mathrm{V}(\mathcal{A})$. Now, note that $\mathrm{K}_{0}$ is functorial for unital $\mathrm{C}^{*}$-algebra homomorphisms in a natural way: if $\phi\colon\mathcal{A}\to B$ is a unital ring homomorphism, and $e\in\mathrm{M}_{\infty}(\mathcal{A})$ is an idempotent, then $\phi(e)$ is an idempotent in $\mathrm{M}_{\infty}(\mathcal{B})$, and the map $e]\mapsto[\phi(e)]$ is well-defined as a map $\mathrm{V}(\mathcal{A})\to\mathrm{V}(\mathcal{B})$. Hence it induces a map $\phi_{*}\colon\mathrm{K}_{0}(\mathcal{A})\to\mathrm{K}_{0}(\mathcal{B})$ by the universal properties of this group
Let $\mathcal{A}$ be a not-necessarily unital $\mathbb{C}$-algebra, and let $\tilde{A}$ be its unitisation, which is equipped with a canonical unital $\mathbb{C}$-algebra homomorphism $\phi\colon\tilde{\mathcal{A}}\to\mathbb{C}$ with kernel $\mathcal{A}$. Then the group $\mathrm{K}_{0}(\mathcal{A})$ is defined by
$\mathrm{K}_{0}(\mathcal{A})\overset{\mathrm{def}}{=}\operatorname{Ker}\phi_{*}\colon\mathrm{K}_{0}(\tilde{\mathcal{A}})\to\mathrm{K}_{0}(\mathbb{C}).$
A local Banach algebra is a normed algebra $\mathcal{A}$ which is closed under holomorphic functional calculus (i.e. if $x\in\mathcal{A}$ and $f$ is an analytic function on a neighborhood of the spectrum of $x$ in the completion of $\mathcal{A}$, with $f(0)=0$ if $\mathcal{A}$ is nonunital, then $f(x)\in\mathcal{A}$.) For technical reasons we will also require that all matrix algebras over $A$ have the same property. If $\mathcal{A}$ is a *-algebra, it will be called a local Banach *-algebra; if the norm is a pre-$\mathrm{C}^{*}$-norm, $\mathcal{A}$ will be called a local $\mathrm{C}^{*}$-algebra.
@@
Example 2.11
If $\mathcal{A}=\mathrm{C}(X)$ is the algebra of a compact Hausdorff space $X$, then by [GVF01, Corollary 3.21 p. 101] we have an isomorphism $\mathrm{K}_{0}(\mathrm{C}(X))=\mathrm{K}^{0}(X)$ where $\mathrm{K}^{0}(X)$ is the topological $\mathrm{K}$-theory group defined by vector bundles.
@@ 3 Banach Fredholm modules and Banach $\mathrm{K}$-homology
In this section, we introduce a Banach space variant of the theory of $\mathrm{K}$-homology, relying on the notion of Fredholm module. Our aim is to replace Hilbert spaces by Banach spaces. Note that the classical notion of Fredholm module admits several variations in the literature (compare the references [HiR00, Definition 8.1.1 p. 199], [Con94, Definition 1 p. 293] and [CGIS14, Definition 2.2]). After the first version of this work appeared, we realized that Lafforgue had had implicitly introduced a similar notion in [Laf02, Definition 1.2.1 p. 16], although with a different purpose, namely in connection with the Baum-Connes conjecture (see [GJV19] for a survey). We begin with the following definition.
@@
Let $\mathscr{B}$ be a class of Banach spaces containing the zero space stable under isomorphisms and finite sums and let $\mathcal{A}$ be an algebra. The Banach $\mathrm{K}$-homology group $\mathrm{K}^{1}(\mathcal{A},\mathscr{B})$ is the abelian group with one generator for each isomorphic equivalence class $(X,\pi,F)]$ of Banach Fredholm modules over $\mathcal{A}$ on any $X$ of $\mathscr{B}$ subject only to the relations:
@@
For any $1\leqslant p<\infty$, it is natural to consider the class $\mathscr{L}^{p}$ of $\mathrm{L}^{p}$-spaces and the class $\mathscr{L}^{p}_{\mathrm{nc}}$ of noncommutative $\mathrm{L}^{p}$-spaces. We could also consider the class $\mathscr{S}\mathscr{L}_{p}$ of subspaces of $\mathrm{L}^{p}$-spaces and the class $\mathscr{S}\mathscr{Q}\mathscr{L}_{p}$ of subspaces of quotients of $\mathrm{L}^{p}$-spaces.
@@
(3.5) $\mathop{}\mathopen{}\mathrm{d}x\overset{\mathrm{def}}{=}[F,x]$
(note that $\mathop{}\mathopen{}\mathrm{d}x\overset{\mathrm{def}}{=}\mathrm{i}[F,x]$ ou $\mathop{}\mathopen{}\mathrm{d}x\overset{\mathrm{def}}{=}2\pi\mathrm{i}[F,x]$ is better for some topics but we do not need it in this paper). A simple computation shows that
(3.6) $\mathop{}\mathopen{}\mathrm{d}(ab)=a\mathop{}\mathopen{}\mathrm{d}b+\mathop{}\mathopen{}\mathrm{d}a\cdot b\quad\text{and}\quad F\mathop{}\mathopen{}\mathrm{d}a\sim-\mathop{}\mathopen{}\mathrm{d}aF.$

### v4

@@ Abstract
The aim of this paper is to bridge noncommutative geometry with classical harmonic analysis on Banach spaces, focusing primarily on both classical and noncommutative $\mathrm{L}^{p}$ spaces. Introducing a notion of Banach Fredholm module, we define new abelian groups, $\mathrm{K}^{0}(\mathcal{A},\mathscr{B})$ and $\mathrm{K}^{1}(\mathcal{A},\mathscr{B})$, of $\mathrm{K}$-homology associated with an algebra $\mathcal{A}$ and a suitable class $\mathscr{B}$ of Banach spaces, such as the class of $\mathrm{L}^{p}$-spaces. We establish index pairings of these groups with the $\mathrm{K}$-theory groups of the algebra $\mathcal{A}$. Subsequently, by considering (noncommutative) Hardy spaces, we uncover the natural emergence of Hilbert transforms, leading to Banach Fredholm modules and culminating in new index theorems. Moreover, by associating each reasonable sub-Markovian semigroup of operators with a ¡¡Banach noncommutative manifold¿¿, we explain how this leads to (possibly kernel-degenerate) Banach Fredholm modules, thereby revealing the role of vectorial Riesz transforms in this context. Overall, our approach significantly integrates the analysis of operators on $\mathrm{L}^{p}$-spaces into the expansive framework of noncommutative geometry, offering new perspectives.
@@ 1 Introduction
A fundamental concept in this framework is that of a Fredholm module. If $\mathcal{A}$ is unital, a Fredholm module over $\mathcal{A}$ is a bounded operator $F\colon H\to H$ satisfying the three relations $^{2}=\mathrm{Id}_{H}, F=F$$F^{*}$ and $F,\pi(a)]=0$ for any $a\in\mathcal{A}$ modulo a compact operator. The last equality is a form of ¡¡pseudolocality¿¿. This notion of Fredholm module is a ¡¡normalized¿¿ variant of the notion of Atiyah cycle1110. An Atiyah cycle on $\mathcal{X}$ is a Fredholm operator $T\colon H_{1}\to H_{2}$ in Hilbert spaces $H_{1}$ and $H_{2}$ equipped with the structure of *-modules over the algebra $\mathrm{C}(\mathcal{X})$ such that the commutator $[T,f]$ is a compact operator for any function $f\in\mathrm{C}(\mathcal{X})$. By Kasparov’s lemma (Higson and Roe 2000), if we assume that the module structures are nondegenerate (i.e., $\mathrm{C}(\mathcal{X})H_{1,2}=H_{1,2}$), then condition (10.6) is equivalent to pseudolocality: the operator $fDg$ is compact for any continuous functions $f$, $g$ on $X$ with disjoint supports. (or generalized elliptic operator) introduced in the paper [Ati70] (see also [BDF77, Definition 7.6 p. 317], [Dou86, Definition 6 p. 165] and [NSSS06, Definition 10.9 p. 274]). If $D$ is the Dirac operator on a compact Riemannian spin manifold $M$, the Fredholm module $\operatorname{\mathrm{sgn}}D$ canonically associated to $M$ encodes the conformal structure of the manifold, as discussed in [Bar07, Theorem 3.1 p. 388]. These operators enable the definition of $\mathrm{K}$-homology groups $\mathrm{K}^{0}(\mathcal{A})$ and $\mathrm{K}^{1}(\mathcal{A})$, whose elements are equivalence classes of Fredholm modules, defined under appropriate notions of unitary equivalence and homotopy equivalence, and a natural operation of addition. The distinction between the two groups lies in the fact that, in the case of $\mathrm{K}^{0}(\mathcal{A})$, only even Fredholm modules are considered, provided an appropriate grading is introduced. These groups are linked to the $\mathrm{K}$-theory groups $\mathrm{K}_{0}(\mathcal{A})$ and $\mathrm{K}_{1}(\mathcal{A})$ of the algebra $\mathcal{A}$ through pairings
@@
of $a$. In the case $F=\mathrm{i}\mathcal{H}$ and $a=f\in\mathrm{C}_{0}^{\infty}(\mathbb{R})$, we have
(1.6) $\big([F,M_{f}]g\big)(x)=\operatorname{p.v.}\int_{-\infty}^{\infty}\frac{f(x)-f(y)}{x-y}g(y)\mathop{}\mathopen{}\mathrm{d}y.$
@@
by integrating over the boundary of a larger bisector $\Sigma^{\mathrm{bi}}_{\nu}$ using the resolvent operator $R(z,D)\overset{\mathrm{def}}{=}(z-D)^{-1}$, where $\omega<\nu<\theta$. Note that the boundedness of such a functional calculus is not free, contrary to the case of the functional calculus of a selfadjoint operator. We refer to our paper [Arh24c] for concrete examples of operator with such a bounded functional calculus, where we use a notion of curvature for obtaining it. Using the function $\operatorname{\mathrm{sgn}}$ defined by $\operatorname{\mathrm{sgn}}(z)\overset{\mathrm{def}}{=}1_{\Sigma_{\theta}}(z)-1_{-\Sigma_{\theta}}(z)$, we show as the hilbertian case that $\operatorname{\mathrm{sign}}D\colon X\to X$ is a Fredholm module. This notion of functional calculus was popularized in the paper [AKM06], which contains a (second) solution to famous Kato’s square root problem solved in [AHLMT02] and in [AKM06] (see also [HLM02] and [Tch01] and the Bourbaki seminar [Mey03]).
@@ 2 Preliminaries
where $(\varepsilon_{k})_{k\geqslant 1}$ is a sequence of independent Rademacher variables on some probability space $\Omega$. As explained in [HvNVW18, Remark 8.1.2 p. 165], this property is independent of $p$. By [HvNVW18, Theorem 8.1.3 p. 166], an $R$-bounded subset of operators is bounded. Finally, a singleton $\{T\}$ is automatically $R$-bounded by [HvNVW18, Example 8.1.7 p. 170]. Actually, the notion of $R$-boundedness can be viewed as a natural and powerful substitute for the usual notion of boundedness, particularly useful when extending results that hold on Hilbert spaces to the more general setting of Banach spaces.
@@
Let $\mathrm{H}^{\infty}_{0}(\Sigma_{\theta})$ be the subalgebra of bounded analytic functions $f\colon\Sigma_{\theta}\to\mathbb{C}$ for which there exist $s,c>0$ such that
@@
The following result combines [NeV17, Proposition 2.3 p. 6] and [HvNVW18, Proposition 10.6.2 p. 448]. It establishes a connection between sectorial and bisectorial operators and their functional calculi.
@@
Let $\omega\in(0,\frac{\pi}{2})$.
  1. 1.
    If $A$ is a bisectorial operator of type $\omega$ then the operator $A^{2}$ is sectorial of type $2\omega$.
  2. 2.
    Suppose that $A$ is an $R$-bisectorial operator on a Banach space $X$ of finite cotype. Then $A^{2}$ is $R$-sectorial and the operator $A$ admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\omega}^{\mathrm{bi}})$-calculus if and only if $A^{2}$ admits a bounded $\mathrm{H}^{\infty}(\Sigma_{2\omega}^{+})$-calculus.
@@
We refer to the books [Bl98], [CMR07], [Eme24], [GVF01], [RLL00] and [WeO93] for more information and to [BaS11] for a short introduction. We briefly recall the algebraic construction of the $\mathrm{K}_{0}$-group, which applies to Banach algebras as well as to more general rings.
$\mathrm{K}_{0}$-group
Let $\mathcal{A}$ be a unital ring. We denote by $\mathrm{M}_{\infty}(\mathcal{A})$ the algebraic direct limit of the algebras $\mathrm{M}_{n}(\mathcal{A})$ under the embeddings $x\mapsto\operatorname{diag}(x,0)$. The direct limit $\mathrm{M}_{\infty}(\mathcal{A})$ can be thought of as the algebra of infinite-dimensional matrice with entries in $\mathcal{A}$, all but finitely many of which are zero. Two idempotents $e,f\in\mathrm{M}_{\infty}(\mathcal{A})$ are said to be (Murray–von Neumann) equivalent if there exist elements $v,w\in\mathrm{M}_{\infty}(\mathcal{A})$ such that $e=vw$ and $f=wv$. The equivalence class of an idempotent $e$ is denoted by $[e]$ and we introduce the set $\mathrm{V}(\mathcal{A})\overset{\mathrm{def}}{=}\{[e]:e\in\mathrm{M}_{\infty}(\mathcal{A}),e^{2}=e\}$ of all equivalence classes of idempotents in $\mathrm{M}_{\infty}(\mathcal{A})$.
According to [CMR07, pp. 3-4], the set $\mathrm{V}(\mathcal{A})$ has a canonical structure of abelian semigroup, with the addition operation defined by the formula $e]+[f]\overset{\mathrm{def}}{=}[\operatorname{diag}(e,f)]$. The identity of the semigroup $\mathrm{V}(\mathcal{A})$ is given by $[0]$, where 0 is the zero idempotent. The group $\mathrm{K}_{0}(\mathcal{A})$ of the unital ring $\mathcal{A}$ is the Grothendieck group of the abelian semigroup $\mathrm{V}(\mathcal{A})$. In particular, each element of $\mathrm{K}_{0}(\mathcal{A})$ identifies with a formal difference [e][f][e]-[f]. Two such formal differences $e_{1}]-[f_{1}]$ and $e_{2}]-[f_{2}]$ are equal in $\mathrm{K}_{0}(\mathcal{A})$ if there exists $g\in\mathrm{V}(\mathcal{A})$ such that $e_{1}]+[f_{2}]+[g]=[f_{1}]+[e_{2}]+[g]$.
@@
Recall that by [CMR07, p. 4 and p. 9], $\mathrm{K}_{0}$ is functorial in a natural way: if $\phi\colon\mathcal{A}\to\mathcal{B}$ is a unital ring homomorphism, and $e\in\mathrm{M}_{\infty}(\mathcal{A})$ is an idempotent, then $\phi(e)$ is an idempotent in $\mathrm{M}_{\infty}(\mathcal{B})$, and the map $\mathrm{V}(\mathcal{A})\to\mathrm{V}(\mathcal{B})$, $e]\mapsto[\phi(e)]$ is well-defined. Hence it induces a group homomorphism $\phi_{*}\colon\mathrm{K}_{0}(\mathcal{A})\to\mathrm{K}_{0}(\mathcal{B})$ by the universal property of the Grothendieck group.
Example 2.11
If $\mathcal{A}=\mathrm{C}(\mathcal{X})$ is the algebra of continuous functions on a second countable compact Hausdorff topological space $\mathcal{X}$, then by [Eme24, Proposition 8.1.7 p. 305] we have an isomorphism $\mathrm{K}_{0}(\mathrm{C}(\mathcal{X}))\cong\mathrm{K}^{0}(\mathcal{X})$ where $\mathrm{K}^{0}(\mathcal{X})$ is the topological $\mathrm{K}$-theory group defined by vector bundles.
If $\mathcal{A}$ is a (not necessarily unital) ring, let $\tilde{\mathcal{A}}=\mathcal{A}\oplus\mathbb{Z}$ be its unitization ring, with the canonical unital homomorphism $\varepsilon\colon\tilde{\mathcal{A}}\to\mathbb{Z}$, $(a,\lambda)\mapsto\lambda$ and kernel $\mathcal{A}$. Following [CMR07, p. 11], we set $\mathrm{K}_{0}(\mathcal{A})\overset{\mathrm{def}}{=}\operatorname{Ker}\varepsilon_{*}$ where $\varepsilon_{*}\colon\mathrm{K}_{0}(\tilde{\mathcal{A}})\to\mathrm{K}_{0}(\mathbb{Z})\approx\mathbb{Z}$.
$\mathrm{K}_{1}$-group
@@ 3 Banach Fredholm modules and Banach $\mathrm{K}$-homology
In this section, we introduce a Banach space variant of the theory of $\mathrm{K}$-homology, relying on the notion of Fredholm module. Our aim is to replace Hilbert spaces by Banach spaces. Note that the classical notion of Fredholm module admits several variations in the literature (compare the references [HiR00, Definition 8.1.1 p. 199], [Con94, Definition 1 p. 293] and [CGIS14, Definition 2.2]). After the first version of this work appeared, we realized that Lafforgue had implicitly introduced a similar notion in [Laf02, Definition 1.2.1 p. 16], although with a different purpose, namely in connection with the Baum-Connes conjecture (see [GJV19] for a survey). We begin with the following definition.
@@
Let $\mathscr{B}$ be a class of Banach spaces containing the zero space stable under finite sums and let $\mathcal{A}$ be an algebra. The Banach $\mathrm{K}$-homology group $\mathrm{K}^{1}(\mathcal{A},\mathscr{B})$ is the abelian group with one generator for each isomorphic equivalence class $(X,\pi,F)]$ of Banach Fredholm modules over $\mathcal{A}$ on any $X$ of $\mathscr{B}$ subject only to the relations:
@@
For any $1\leqslant p<\infty$, it is natural to consider the class $\mathscr{L}^{p}$ of spaces isomorphic to $\mathrm{L}^{p}$-spaces and the class $\mathscr{L}^{p}_{\mathrm{nc}}$ of spaces isomorphic to noncommutative $\mathrm{L}^{p}$-spaces. We could also consider the class $\mathscr{S}\mathscr{L}_{p}$ of saces isomorphoc to subspaces of $\mathrm{L}^{p}$-spaces and the class $\mathscr{S}\mathscr{Q}\mathscr{L}_{p}$ of spaces isomorphic to subspaces of quotients of $\mathrm{L}^{p}$-spaces.
@@
[…]